Conditional Distribution
Question.Of employed u.s adults age 25 or older, 90.3% have completed high school, while 30.8% have completed college. for H = is completed high school, C = completed college, and assuming that one must complete high school before completing college, find following 
a) P(H and C)
b)P(C/H)
c) P(C'/H)
d) formula and solution
Thank you

Answer by stanbon(75887)

Conditional Distribution
Question.Of employed u.s adults age 25 or older, 90.3% have completed high school, while 30.8% have completed college. for H = is completed high school, C = completed college, and assuming that one must complete high school before completing college, find following
==================
P(H) = 0.903 ; P(C) = 0.308
a) P(H and C) = 0.308
----------------------------
b) P(C|H) = P(C and H)/P(H) = 0.308/0.903
----------------------------
c) P(C'|H) = P(C' and H) = (0.692*0.903 = 0.625
-------------------------------
d) formula and solution ?????
Cheers,
Stan H.
===========